Form an AM DSBFC envelope with a 40V positive maximum voltage and a 10V positive minimum voltage, determine (a) the unmodulated carrier amplitude, (b) the peak change in amplitude of the modulated wave, and (c) coefficient of modulation and percent modulation.



1
Expert's answer
2021-12-09T09:22:40-0500

a)

"E_c=(V_{max}+V_{min})\/2=(40+10)\/2=25" Vp


b)

"E_m=(V_{max}-V_{min})\/2=(40-10)\/2=15" Vp


c)

coefficient of modulation:

"m=E_m\/E_c=15\/25=0.6"


percent modulation:

"M=100m=0.6\cdot100=60\%"